# Resonant or asymmetric: The status of sub-GeV dark matter

### v1

@@ 1 Introduction
The most plausible explanation for the wealth of astrophysical and cosmological evidence for dark matter (DM) is the existence of new elementary particles. Without additional assumptions, however, the allowed mass range for such particles spans almost 50 orders of magnitude, with the lower bound of around $10^{-30}\,\mathrm{GeV}$ set by quantum mechanics (the De Broglie wavelength must be smaller than the size of known astrophysical objects [1, 2]) and the upper bound of around $10^{19}\,\mathrm{GeV}$ set by gravity (the mass of a fundamental particle must be smaller than the Planck mass). Even larger DM masses are possible when considering composite states. This enormous window can be narrowed down significantly, if we consider DM particles that obtain their relic abundance in the early universe by freezing-out from the SM thermal bath. In this case, a lower bound on the DM mass of around 10 MeV is imposed by cosmology, specifically constraints on the number of relativistic degrees of freedom from Big Bang Nucleosynthesis (BBN) and the Cosmic Microwave Background (CMB) [3, 4], and an upper bound of around 100 TeV is imposed by the unitarity requirement that the annihilation cross section of a particle cannot be arbitrarily large [5]. Over the past few decades, the central part of this range, from a few GeV to a few TeV, has been the target of a large number of direct and indirect detection experiments and collider searches, and their null results have strongly constrained the corresponding DM models [6, 7]. For many models, the leading constraints come from direct detection experiments searching for nuclear recoils, which however rapidly lose sensitivity for DM particles below the GeV scale, because the typical kinetic energy of such particles lies below the detector threshold.
@@
Nevertheless, models of sub-GeV DM face a key challenge, which is that they need to satisfy strong constraints on the DM annihilation cross section from the CMB [13, 14] and, as recently pointed out in ref. [15], from searches for X-ray emission. There are essentially three ways to evade these constraints: by considering a strongly velocity-dependent annihilation cross section arising for example from $p$-wave suppression [16] or resonant enhancement [17, 18], by allowing for a particle-antiparticle asymmetry suppressing annihilation rates in the present universe [19], or by considering DM particles that only constitute a fraction $f$ of the total DM density, such that annihilation signals are suppressed proportional to $f^{2}$ even in the absence of an asymmetry. In the present work we consider all three possibilities simultaneously and determine which ones are favoured or disfavoured by data. [^1] Another possibility would be to consider alternative ways for DM particles to obtain their relic abundance, such as secluded or forbidden annihilations, or non-thermal production via the freeze-in mechanism. These possibilities imply however considerably smaller couplings that are much harder to probe experimentally and therefore outside the scope of the present work.
@@ 2.1 Decays, annihilations and self-interactions
In principle, DM particles could also annihilate into dark photon pairs, but for the spectrum that we consider ($m_{A^{\prime}}>2m_{\text{DM}}$) this decay mode is strongly suppressed and plays a negligible role in the calculation of the DM relic abundance.
@@ 3.1.3 Exotic energy injection
Finally, we note that for the smallest DM masses that we consider, there may be relevant constraints from DM annihilations after the end of BBN, leading to the photodisintegration of deuterium and helium-3 [3]. However, it was argued in ref. [17] that this constraint is only relevant for parameter points with very large resonant enhancement (corresponding to $\epsilon_{R}<10^{-3}$). These parameter regions are not the focus of the present study, and we therefore leave a detailed study of constraints from photodisintegration to future work [62].
@@ 3.2 Astrophysical constraints
Assuming that DM is not fully asymmetric, it may annihilate to SM particles, producing detectable gamma-ray or X-ray signals. Annihilation directly to photons would lead to a distinctive monochromatic line signal at the DM mass, while annihilation to quarks, leptons and mesons can result in a broad photon spectrum resulting from the subsequent particle fragmentation and decay. Searches for both signals can set strong limits on sub-GeV DM (for example refs. [63, 64]). However, gamma-ray searches for sub-GeV dark matter are limited by the so-called “MeV gap” between keV-range experiments like SPI aboard the INTEGRAL satellite [65], and Fermi-LAT at GeV energies [66]. Numerous experiments have been proposed to fill in this gap in coverage; for a detailed review, see ref. [67].
@@
We note that ref. [77] recently updated the results of ref. [15] using a more realistic treatment of cosmic ray propagation including spatial diffusion and reacceleration. While the resulting limits can be substantially stronger, especially at low mass, they depend on the model and parameters used to describe cosmic ray transport, which carry large uncertainties. We therefore conservatively use the earlier result [15] in the present work.
@@
Merging galaxy cluster 1E0657-56, otherwise known as the Bullet Cluster, serves as one of the best test beds for DM self-interactions. It comprises a subcluster that passed through the main cluster’s central region along the plane of the sky. X-ray and lensing observations have identified gas and mass distributions with well-separated peaks, implying that feebly interacting DM dominates the clusters’ masses [78, 79, 80, 81]. The presence of DM self-interactions would lead to the subcluster DM component experiencing friction as it passes through the main cluster. This would result in decelerating the DM component of the subcluster, such that it lags behind the collisionless galaxies, and cause the subcluster to lose part of its DM mass.
@@ 3.3 Accelerator experiments
As discussed in section 2.1, we focus our analysis on parameter regions where the dark photon decays almost exclusively into pairs of DM particles. Constraints on this scenario can come from two types of experiments: searches for missing energy and searches for the scattering of DM particles produced in the dark photon decays. Figure 5 shows the 90 % C.L. exclusion bounds from LSND [91, 92], MiniBooNE electron and nucleon scattering [93], NA64 [94] and BaBar [95] for a complex scalar DM candidate with an example benchmark choice of model parameters. These exclusion bounds are shown from both previous literature results (dashed) and from using the interpolations and scaling techniques implemented in this work (solid). The rest of this section describes these experiments in detail.
@@
Beam dump experiments such as LSND [91, 92] and MiniBooNE [93] aim to produce a relativistic flux of DM particles from interactions between a proton beam and nucleons in a dense target. Neutral mesons such as $\pi^{0}$ and $\eta$ are produced in these interactions, which then decay producing DM via the chain $\pi^{0},\eta\to\gamma+A^{\prime}$, $A^{\prime}\to\chi\bar{\chi}$ with either $\chi=\psi$ for fermionic DM or $\chi=\Phi$ for scalar DM. The DM particles can then be detected in a downstream detector through DM-nucleon or DM-electron scattering, analogous to underground DM direct detection experiments searching for Galactic DM. The number of expected signal events in the detector is proportional to model parameters $g_{\text{DM}}$ and $\kappa$, and branching ratios $\text{BR}_{X\to A^{\prime}\gamma}$ and $\text{BR}_{A^{\prime}\to\chi\bar{\chi})}$:
@@
  • The visible branching fraction satisfies the constraint $\kappa^{2}\text{BR}_{A^{\prime}\to SM}<10^{-7}$, such that prompt decays into SM final states are out of reach of the leading experiments such as BaBar [101] and NA48/2 [102].
@@ 3.4 Direct detection
The leading constraint on sub-GeV DM from nuclear recoils stems from CRESST-III [115], which we implement using DDCalc [30]. However, substantially stronger constraints on the DM-proton scattering cross section can be obtained by exploiting the Migdal effect, which has been implemented in obscura for liquid noble gas detectors following refs. [116, 117, 118]. We use this implementation and the electron recoil data discussed above to calculate likelihoods for DarkSide50, XENON1T and PandaX-4T, finding that PandaX-4T gives the strongest constraint.
@@ 4.1 Fermionic dark matter
  • Finally, in the bottom-right panel, we can see that for increasing $g_{\text{DM}}$ the allowed range of $\kappa$ shrinks towards the upper bound. This is because the relic density requirement is violated if both couplings $g_{\text{DM}}$ and $\kappa$ become small simultaneously (see figure 2).
@@
It is furthermore instructive to plot various observables as a function of the DM mass, as shown in figure 8. We can see that in the bulk of the parameter space the DM relic abundance is well below the observed value, such that indirect detection signals (which scale proportional to $f_{\text{DM}}^{2}$ for symmetric DM) and direct detection signals (which scale proportional to $f_{\text{DM}}$) can be strongly suppressed. Nevertheless, the observed DM relic abundance can be saturated across the entire DM mass range. Both direct and indirect detection experiments place relevant constraints on the parameter space, which can however be evaded by many orders of magnitude when considering sub-dominant dark matter components. However, it turns out that the predicted DM-electron scattering cross section is below current exclusion limits in the entire allowed parameter space.
@@ 4.2 Scalar dark matter
Figure 15: Allowed parameter regions for symmetric scalar dark matter with $\Omega_{\text{DM}}h^{2}\approx 0.12$. Gray lines indicate the allowed parameter regions for a sub-dominant DM component and the gray star indicates the corresponding best-fit point. See caption of figure 7 for details on the various line styles and symbols.
@@ 5.1 Sensitivity projections
In figure 18 we compare the allowed parameter regions of the two preferred models, namely asymmetric fermionic DM and symmetric scalar DM, with the projected sensitivities of near-future experiments. For accelerator experiments we consider the single-photon search of Belle II with 20 $\mathrm{fb^{-1}}$ [128] and the missing energy search of LDMX with $10^{16}$ electrons from an 8 GeV beam [129]. For DM-nucleus scattering we consider SuperCDMS [130], and the recently proposed DELight experiment [131], which plans to use superfluid helium to reach lower DM masses. For DM-electron scattering we show the projections from DAMIC and SENSEI (both taken from ref. [114]).
@@ E.2 Updates to DarkBit, CosmoBit and ColliderBit
In DarkBit, many capabilities are supplemented with module functions for the sub-GeV models, including those for computing the process catalog, TH_ProcessCatalog_SubGeVDM_fermion and TH_ProcessCatalog_SubGeVDM_scalar, those that calculate the the direct detection couplings, DD_couplings_SubGeVDM_fermion and DD_couplings_SubGeVDM_scalar, and those storing the DM properties, DarkMatter_ID_SubGeVDM_fermion and DarkMatterConj_ID_SubGeVDM_fermion, for fermion DM, and DarkMatter_ID_SubGeVDM_scalar, and DarkMatterConj_ID_SubGeVDM_scalar for scalar DM.

### v2

@@ 1 Introduction
The most plausible explanation for the wealth of astrophysical and cosmological evidence for dark matter (DM) is the existence of new elementary particles. Without additional assumptions, however, the allowed mass range for such particles spans almost 50 orders of magnitude, with the lower bound of around $10^{-30}\,\mathrm{GeV}$ set by quantum mechanics (the De Broglie wavelength must be smaller than both the size of known astrophysical objects [1] and the correlation length of structures in the Lyman alpha forest [2, 3]) and the upper bound of around $10^{19}\,\mathrm{GeV}$ set by gravity (the mass of a fundamental particle must be smaller than the Planck mass). Even larger DM masses are possible when considering composite states. This enormous window can be narrowed down significantly, if we consider DM particles that obtain their relic abundance in the early universe by freezing-out from the SM thermal bath. In this case, a lower bound on the DM mass of around 10 MeV is imposed by cosmology, specifically constraints on the number of relativistic degrees of freedom from Big Bang Nucleosynthesis (BBN) and the Cosmic Microwave Background (CMB) [4, 5], and an upper bound of around 100 TeV is imposed by the unitarity requirement that the annihilation cross section of a particle cannot be arbitrarily large [6]. Over the past few decades, the central part of this range, from a few GeV to a few TeV, has been the target of a large number of direct and indirect detection experiments and collider searches, and their null results have strongly constrained the corresponding DM models [7, 8]. For many models, the leading constraints come from direct detection experiments searching for nuclear recoils, which however rapidly lose sensitivity for DM particles below the GeV scale, because the typical kinetic energy of such particles lies below the detector threshold.
@@
Nevertheless, models of sub-GeV DM face a key challenge, which is that they need to satisfy strong constraints on the DM annihilation cross section from the CMB [14, 15] and, as recently pointed out in ref. [16], from searches for X-ray emission. There are essentially three ways to evade these constraints: by considering a strongly velocity-dependent annihilation cross section arising for example from $p$-wave suppression [17] or resonant enhancement [18, 19, 20], by allowing for a particle-antiparticle asymmetry suppressing annihilation rates in the present universe [21], or by considering DM particles that only constitute a fraction $f$ of the total DM density, such that annihilation signals are suppressed proportional to $f^{2}$ even in the absence of an asymmetry. In the present work we consider all three possibilities simultaneously and determine which ones are favoured or disfavoured by data. [^1] Another possibility would be to consider alternative ways for DM particles to obtain their relic abundance, such as secluded or forbidden annihilations, or non-thermal production via the freeze-in mechanism. These possibilities typically imply considerably smaller couplings that are much harder to probe experimentally and therefore outside the scope of the present work.
@@ 2.1 Decays, annihilations and self-interactions
In principle, DM particles could also annihilate into dark photon pairs, but for the spectrum that we consider ($m_{A^{\prime}}>2m_{\text{DM}}$) this decay mode is strongly suppressed and plays a negligible role in the calculation of the DM relic abundance. We also note that ref. [47] recently showed that higher-order corrections to the annihilation cross section may be non-negligible for large values of $g_{\mathrm{DM}}$. Given other large theory uncertainties that enter the relic density calculation (see below), these corrections are not included in the present work.
@@ 3.1.3 Exotic energy injection
Finally, we note that for the smallest DM masses that we consider, there may be relevant constraints from DM annihilations after the end of BBN, leading to the photodisintegration of deuterium and helium-3 [4]. However, it was argued in ref. [18] that this constraint is only relevant for parameter points with very large resonant enhancement (corresponding to $\epsilon_{R}<10^{-3}$[65]. These parameter regions are not the focus of the present study, and we therefore leave a detailed study of constraints from photodisintegration to future work.
@@ 3.2 Astrophysical constraints
Assuming that DM is not fully asymmetric, it may annihilate to SM particles, producing detectable gamma-ray or X-ray signals. Annihilation directly to photons would lead to a distinctive monochromatic line signal at the DM mass, while annihilation to quarks, leptons and mesons can result in a broad photon spectrum resulting from the subsequent particle fragmentation and decay. However, gamma-ray searches for sub-GeV dark matter are limited by the so-called “MeV gap” between keV-range experiments like SPI aboard the INTEGRAL satellite [66], and Fermi-LAT at GeV energies [67]. Numerous experiments have been proposed to fill in this gap in coverage; see ref. [68] for a detailed review. However, for the models that we consider in this work, in which DM particles annihilate primarily into charged leptons and mesons, while neutral mesons – and hence photons – are sub-dominant, MeV gamma-ray telescopes are not expected to improve on the CMB constraints discussed above [69].
@@
We note that ref. [79] recently updated the results of ref. [16] using a more realistic treatment of cosmic ray propagation including spatial diffusion and reacceleration, and ref. [80] studied additional constraints obtained from INTEGRAL measurements of the 511 keV line. While the resulting limits can be substantially stronger, especially at low mass, they depend on the model and parameters used to describe cosmic ray transport, which carry large uncertainties. We therefore conservatively use the earlier result [16] in the present work. These constraints are still stronger than the ones obtained from Voyager 1 [81].
@@
Merging galaxy cluster 1E0657-56, otherwise known as the Bullet Cluster, serves as one of the best test beds for (velocity-independent) DM self-interactions. It comprises a subcluster that passed through the main cluster’s central region along the plane of the sky. X-ray and lensing observations have identified gas and mass distributions with well-separated peaks, implying that feebly interacting DM dominates the clusters’ masses [82, 83, 84, 85]. The presence of DM self-interactions would lead to the subcluster DM component experiencing friction as it passes through the main cluster. This would result in decelerating the DM component of the subcluster, such that it lags behind the collisionless galaxies, and cause the subcluster to lose part of its DM mass.
@@ 3.3 Accelerator experiments
As discussed in section 2.1, we focus our analysis on parameter regions where the dark photon decays almost exclusively into pairs of DM particles. Constraints on this scenario can come from two types of experiments: searches for missing energy and searches for the scattering of DM particles produced in the dark photon decays. Figure 5 shows the 90 % C.L. exclusion bounds from LSND [95, 96], MiniBooNE electron and nucleon scattering [97], NA64 [98] and BaBar [99] for a complex scalar DM candidate with an example benchmark choice of model parameters. These exclusion bounds are shown from both previous literature results (dashed) and from using the interpolations and scaling techniques implemented in this work (solid). The rest of this section describes these experiments in detail. Note that we do not consider the recent MicroBooNE analysis [100], which employs a convolutional neural network and therefore cannot easily be reinterpreted in our context. For similar reasons, we also do not include the first results from the COHERENT experiment [101].
@@
Beam dump experiments such as LSND [95, 96] and MiniBooNE [97] aim to produce a relativistic flux of DM particles from interactions between a proton beam and nucleons in a dense target. Neutral mesons such as $\pi^{0}$ and $\eta$ are produced in these interactions, which then decay producing DM via the chain $\pi^{0},\eta\to\gamma+A^{\prime}$, $A^{\prime}\to\chi\bar{\chi}$ with either $\chi=\psi$ for fermionic DM or $\chi=\Phi$ for scalar DM. The DM particles can then be detected in a downstream detector through DM-nucleon or DM-electron scattering, analogous to underground DM direct detection experiments searching for Galactic DM. The number of expected signal events in the detector is proportional to model parameters $g_{\text{DM}}$ and $\kappa$, and branching ratios $\text{BR}_{X\to A^{\prime}\gamma}$ and $\text{BR}_{A^{\prime}\to\chi\bar{\chi}}$:
@@
  • The visible branching fraction satisfies the constraint $\kappa^{2}\text{BR}_{A^{\prime}\to SM}<10^{-7}$, such that prompt decays into SM final states are out of reach of the leading experiments such as LHCb [107], BaBar [108] and NA48/2 [109].
@@ 3.4 Direct detection
The leading constraint on sub-GeV DM from nuclear recoils stems from CRESST-III [122], which we implement using DDCalc [32]. However, substantially stronger constraints on the DM-proton scattering cross section can be obtained by exploiting the Migdal effect, which has been implemented in obscura for liquid noble gas detectors following refs. [123, 124, 125]. We use this implementation and the electron recoil data discussed above to calculate likelihoods for DarkSide50, XENON1T and PandaX-4T [126], finding that PandaX-4T gives the strongest constraint.
@@ 4.1 Fermionic dark matter
  • Finally, in the bottom-right panel, we can see that for increasing $g_{\text{DM}}$ the allowed range of $\kappa$ shrinks towards the upper bound. This is because larger values of $g_{\text{DM}}$ correspond to a larger dark photon width, which implies less resonant enhancement even for small $\epsilon_{R}$.
@@
It is furthermore instructive to plot various observables as a function of the DM mass, as shown in figure 8. We can see that in the bulk of the parameter space the DM relic abundance is well below the observed value, such that indirect detection signals (which scale proportional to $f_{\text{DM}}^{2}$ for symmetric DM) and direct detection signals (which scale proportional to $f_{\text{DM}}$) can be strongly suppressed. Nevertheless, the observed DM relic abundance can be saturated across the entire DM mass range. Both direct and indirect detection experiments place relevant constraints on the parameter space, which can however be evaded by many orders of magnitude when considering sub-dominant dark matter components. However, it turns out that the predicted DM-electron scattering cross section is below current exclusion limits in the entire allowed parameter space. We have also checked explicitly that the visible branching ratio of the dark photon into Standard Model particles is sufficiently small to satisfy all experimental constraints.
@@ 4.2 Scalar dark matter
Figure 15: Allowed parameter regions for symmetric scalar dark matter with $\Omega_{\text{DM}}h^{2}\approx 0.12$. See caption of figure 7 for details on the various line styles and symbols.
@@ 5.1 Sensitivity projections
In figure 18 we compare the allowed parameter regions of the two preferred models, namely asymmetric fermionic DM and symmetric scalar DM, with the projected sensitivities of near-future experiments. For accelerator experiments we consider the single-photon search of Belle II with 20 $\mathrm{fb^{-1}}$ [136] and the missing energy search of LDMX with $10^{16}$ electrons from an 8 GeV beam [137]. For DM-nucleus scattering we consider SuperCDMS [138], and the recently proposed DELight experiment [139], which plans to use superfluid helium to reach lower DM masses. [^8] Very similar sensitivities are expected to be achieved by the HeRALD detector within the TESSERACT programme [140]. For DM-electron scattering we show the projections from DAMIC and SENSEI (both taken from ref. [121]).
@@ E.2 Updates to DarkBit, CosmoBit and ColliderBit
In DarkBit, many capabilities are supplemented with module functions for the sub-GeV models, including those for computing the process catalog, TH_ProcessCatalog_SubGeVDM_fermion and TH_ProcessCatalog_SubGeVDM_scalar, those that calculate the direct detection couplings, DD_couplings_SubGeVDM_fermion and DD_couplings_SubGeVDM_scalar, and those storing the DM properties, DarkMatter_ID_SubGeVDM_fermion and DarkMatterConj_ID_SubGeVDM_fermion, for fermion DM, and DarkMatter_ID_SubGeVDM_scalar, and DarkMatterConj_ID_SubGeVDM_scalar for scalar DM.